What potential solution did you obtain? Explain why this is not a solution. show the steps that you would take to solve the following algebraaiclly 5-2/x-6=10-2x/x-6
Answer by mizjoice(22)

5-2/x-6=10-2x/x-6,
(x-6)5-2/x-6*x-6=10-2x/x-6*x-6,
(x-6)5-2=10-2x,
5x-30-2=10-2x
5x-32+2x=10-2x+2x,
7x-32=10,
7x-32+32=10+32,
7x=42
7x/7=42/7
x=6
To check, substitute 6 for the variable x in the original question.
5-2/6-6=10-12/6-6,
5-2/0=-2/0.
This is not a solution because division by a zero value is not possible.